OLD LNP-10 Log Book

Table of contents
No headers

Below is a checklist for the LNP-10 which should be done daily when the compressors are running or whenever possible when LN2 exists in the Cryomech Dewar but the system is off. 

 

 

Cryomech LNP-10 Checklist
Date & Time Norhof LN2 Level (cm)

Drain 

Air 
Comp.    

Air 

Comp.

Oil 

Lvl

Air

Comp.

Oil

Pres.

PSI

Air

Comp. 

Off

Period

Filter

Bank

Pres.

PSI

Water

Filter

 

ADV

Check

Sight

Carbon &

water

Filter

sight glass

Drain

carbon

filter

Dewar

Pres.

PSI

Dewar

LN2

Level

He

Low

Pres.

PSI

He

High

Pres.

PSI

He 

Comp.

Hours.

 

Air

Dryer

Dew
Point

Air

Dryer

water

Level

Notes:_____________________________________

2016/08/11  

  N/A N/A N/A N/A 85 N/A N/A N/A N/A 7 N/A 200/60-65 195/270-275 N/A N/A N/A These are the stats from the Cryomech service, LN2 production was 10.5 L/day (i.e. Dewar full in 3.3 days)
 2016/09/22 12:09    Y, 100 mL  H  20 6.5   85  G  Y  Y  Y  7.8 1/16  70-80  280-285  680  G  Y This was the first readings taken when starting the unit up in ETL

 2016/10/18 10:52

   N  H  N/A  N/A  0  G  N  N  N  10.25  1/2  210  210  799.5  N/A  empty In standby.
 2016/10/27 17:14
   N  H  18  5  85  G  Y  Y  Y  7  E  210  210  799.5  G  E Dewar Flush
 2016/10/28 13:46
   Y 75 mL
 H  18  5  85  G  Y  Y  Y  6.5  1/4  60  290  819.5  G  1/32

 In Use    

 2016/10/29 18:25

   Y 75mL  H  18  5  85  G  Y  Y  Y  6.5  3/4  60  290  848.1  G  2/32  In Use
 2016/10/30 18:07    Y 75 mL  H  18  5  85  G  Y  Y  Y  6.5  1/2  60  290  872.0  G  E  Filled Norhof
 2016/10/31 20:13    Y75mL  H  18  5  85  G  Y  Y  Y  7  7/8  65-75  285-290  898.0  G  some  In Use
2016/11/1 15:30    Y75mL  N/A  N/A  N/A  0  G  N  Y  N  10.25  15/16  210  205  902.5  N/A  E  Shutdown
 2016/11/10 17:42    N  H  19  5  85  G  Y  Y  Y  10  3/16  65-75  285-290  902.6  G  E  Startup
 2016/11/11 16:00    Y75mL H  18  5  85  G  Y  Y  Y  7  5/8  65-75  285-290  924.9  G  SOME  IN USE
 2016/11/12 17:02    Y75ML  H  18  5  85  G  Y  Y  Y  6  3/8  65-75  285-290  950.3  G  SOME  FILLED NORHOFF
 2016/11/13 22:00   Y 75 mL  N/A  N/A  N/A  0  G  Y  Y  Y  7  F  210  205  978.9  G  some  Shutdown. Cryomec LN2 production rate. 10 Nov 6.52 L, 13 Nov+ 4.3 hr , 37.22 L produced (accounting for Norhof extraction). Places LN2 rate at 11.7 L/day. Slightly above what the factory got which could be accounted for inaccurate LN2 level  measurements.
2016/11/24 16:40   N H  19 5 85  G  Y  Y  N  10.25  1/8  210  205  978.9  G  E  Startup
2016/11/25 18:06   Y, very little H  18  5  85  g  y  y  y  6.5  1/2  65-75  285-290  1004.3  g  some  in use
 2016/11/26 20:39   Y very little  H  N/A  N/A  85  g  y  y  y  9  F  210  205  1030.8  g  some  shutdown
2016/12/1 15:00    N  H  21  5  85  G  y  y  y  8  1/16  210  205  1030.8  G  some  startup, filled norhof, performed 1st oil change on air compressor as required by users manual
 2016/12/2 16:53    Y, 10 drops  H 19  5  85  g y y y  7  3/8  65-75  285-290  1056.3  G  some  IN USE
 2016/12/4 15:11    Y 25 ml  H  N/A  N/A  85  G  Y  Y  Y  10.5  7/8  210  205  1086.9 G  some  Shutdown
 2016/12/14 11:18    N  H  20.5  5  85  G  Y  Y  Y  7  1/8  65-75  285-290  1086.9  G  E

 Startup. Higher oil pressure

possibly due to oil not being hot yet

2016/12/15 16:26    Y very little  H  20  5  85  g  y  y  y  8  E  65-75  285-290  1116.2  G  some  filled norhof
 2016/12/17 09:12    Y, none  H  21  5  85  G  Y  Y  Y  7  5/8  65-75  285-290  1156.8  G  some

 in use. adjusted oil pressure to 18.5 PSIG

Spec. is 18-20 PSIG

2016/12/17 17:42    y, none  H  18.5  5  85  G  Y  Y  Y  7.25  7/8  65-75  285-290  1165.4  G  some  Shutdown
2016/12/22 17:03    N  H  N/A  N/A  0  g  N  
N
 
N
 10.5  1/2  210  205  1165.4  G  some  checkup
 2016/12/29 16:33    N  H  18.5  4.5  85  g  n  n  n  7.5  E  210  205  1165.4  g  E

 Dewar empty on arrival, will flush for 1 hour

and then generate LN2

2016/12/30    Y, none  H  18.5  5  85  G  Y  Y  Y  7.0  1/8  65-75  285-290  1182.0  g  E

Dewar pressure was at 5.75. adjusting

pressure seems to make the needle jump

for a short time.

2017/1/3 8:31    Y, none  H  18.5  5  85  G  y  y  y  6.5  E  65-75  285-290  1229.4  G  E  filled norhof
2017/1/4 17:19    Y, none  H  18.5  5  85  G  Y  Y  Y  6.5  1/2  65-75  285-290  1262.0  G  E  
2017/1/5 15:39  27.4  Y, none  H  N/A  N/A  0  G  Y  Y  Y  10.25  F  210  205  1282.1  G  E  Shutdown
 2017/1/12 16:16  19.3  N  H  N/A  N/A  0  G  Y  Y  Y  10.25  3/8  210  205  1282.1  G  E  STARTUP
 2017/1/12 22:40  18.7  Y, NONE  H  N/A  NA/  85  G  Y  Y  Y  6.5  1/2  210  205  1287.8  G  E

 Shutdown due to high temp.

caused by HVAC failure

 2017/1/13 14:20  18.6  N   H  18.5  5  85  G  Y  Y  Y  10  7/16  65-75  285-290  1288.0  G  E  startup
 2017/1/14 13:29  17.3  Y  H  18.5  5  85  G  Y  Y  Y  6.5  13/16  210  205  1311.1  G  E  shutdown
 2017/1/16  15.3  N  N/A  N/A  N/A  0  G  N  N  N  10.25  5/8  210  205  1311.1  G  E  checkup
 2017/1/17 10:46  13.9  N  H  18.5  5  85  G  N  N  N  10.25  1/2  210  205  1311.1  G  E  STARTUP
 2017/1/19 08:00  38.8  Y,NONE  H  18.5  5  85  G  Y  Y  Y  7  1/4  65-75  285-290  1335.8  G  E  filled norhof
 2017/1/21 14:07  28.7  Y, 3ML  H  N/A  N/A  85  G  Y  Y  Y  10.25  F  210  205  1375.2  G  E  shutdown
 2017/1/26 16:08  22.7  N/A  H  N/A  N/A  0  G  N  N  N  10.25  1/2  210  205  1375.2  G  E  CHECKUP
 2017/1/28 15:33  20.7  N/A  H  19  N/A  85  G  N  N  N  10.25  1/4  210  205  1375.2  G  E  startup
 2017/1/31 11:37  18  Y 10ML  H  18.5  5  85  G  Y  Y  Y  10.25  F  210  205  1414.4  G E  SHUTDOWN
 2017/2/2 16:43  40  Y none  H  17  5  85  G  Y  Y  Y  10.25  E  210  205  1414.4  G  E  filled norhof
 2017/2/7 15:20  25.4  y none  H  N/A  N/A  85  G  Y  Y  Y  10.25  F  210  205  1479.8  G  E  shutdown
 2017/2/9 17:00  22.7  N  H  NA  NA  0  G  N  N  N  10.25  3/4  210  205  1479.8  G  E  CHECK
 2017/2/16 15:31  15.3  N  H  18  5  85  G  N  N  N  10.25  1/4  210  205  1479.8  G  E  startup
 2017/2/17 12:42  12.6  y none  H  15  5  85  g  y  y  y  5.5  1/2  65-75  285-290  1500.4  g  e  checkup, suspect low oil pressure due to heat, will monitor and adjust if needed.
 2017/2/18 11:05  42.2  y none  h  16.5  5  85  g  y  y  y  6.25  1/4  65-75  285-290  1522.9  g  e  filled norhof,
 2017/2/19 15:46  28.7  Y NONE  H  16  5  85  G  Y  Y  Y  6.5  3/4  65-75  285-290  1551.5  G  E  checkup
 2017/2/20 17:40  28  y none  h  n/a  n/a  85  g  y  y  y  10.25  F  210  205  1566.8  g  e  shutdown
 2017/2/23 12:23  25.4  N  H  N/A  N/A  0  G  N  N  N  10.25  5/8  210  205  1566.8  G  E  CHECKUP
 2017/3/2 15:45  18.0  N  H  18  5  85  G  N  N  N  10.25  1/8  210  205  1566.8  G  E  startup
 2017/3/4 09:37  15.3  Y, NONE  H  18  5  85  G  Y  Y  Y  6.25  3/4  65-75  285-290  1608.5  G  E  Filling norhof. noticed that temp. did not spike this time.
 2017/3/6  30.7  y none  h  n/a  n/a  0  g  y  y  y  10.25  f  210  205  1650.3  g  e  shutdown
 2017/3/9 17:29  26  N  H  N/A  N/A  0  G  N  N  N  10.25  1/2  210  205 1650.3  G  E  check up
 2017/3/10 11:42  24.7  N  H  N/A  N/A  0  G  N  N  N  10.25  1/2  210  205  1650.3  G  E  CHECK
 2017/3/16 16:31  18  n  h  18  5  85  g  n  n  n  10.25  1/8  210  205  1650.3  g  e  startup
 2017/3/19 18:35  88.5  y none  h  18  5  85  g  y  y  y  10.25  1/4  65-75  285-290  1703.5  g  e  filled norhof
 2017/3/23  26.7  y none  h  18  5  85  g  y  y  y  6  f  65-75  285-290  1749.5  g  e  shutdown
 2017/3/30  18.6  N  H  18  5  85  G  N  N  N  10.25  3/8  210  205  1749.5  G  some  startup
 2017/4/6 16:49  full  y 100ml  H  18  5  85  g  y  y  y  6.5  1/4  65-75  285-290  1810.5  g  some  filled norhof
 2017/4/8 17:23  29.4  y 75 ml  h  n/a  n/a  0  g  y  y  y  10.25  F  210  205

 1853.9

 g  some  shutdown
 2017/4/13 16:52  22.7  N  H  18  5  85  G  N  N  N  10.25  3/8  210  205  1853.9  G  EMPTY  startup
 2017/4/20 16:05  F  Y 125ML  H  18  5  85  G  Y  Y  Y  10  3/8  65-75  285-290  1904.7  g  empty  filled norhof
 2017/4/21 14:08  30.7  Y 75 ML  H  18  5  85  G  Y  Y  Y  6.5  3/4  65-75  285-290  1926.5  g  some  checkup
 2017/4/22 14:25  29.4  y 50ml  h  18  n/a  85  g  y  y  y  10.25  F  210  205  1942.1  g  some  shutdown
 2017/4/27 17:19  23.3  N  H  18  5  85  G  N  N  N  10.5  1/2  65-75  285-290  1942.1  G  some  startup
 2017/4/28 15:54  22  Y very little H  18  N/A  85  G  Y  Y  Y  6.5  7/8  210  205  1964.7  G  some  Shutdown
 2017/5/4 15:58  15.3  N  H  0  N/A  0  G  N  N  N  10.25  1/8  210  205  1964.7  G  SOME  Sartup
 2017/5/7 12:17  11.9  y 50ml  h  18  5  85  g  y  y  y  10.25  F  210  205  2005.8  g  some  filling norhof
 2017/5/8 08:14  30.7  Y 75ML  H  18  5  85  G  Y  Y  Y  6.75  1/2  65-75  285-290  2025.8  G  some  checkup
 2017/5/11 17:10  24.7  y 75ml  h  18  5  85  g  y  y  y  6.5  7/8+  210  205  2061.7  g  some  shutdown
 2017/5/19 17:34  18  N  H  18  N/A  85  G  N  N  N  10.25  1/8  210  205  2061.7  G  empty  startup
 2017/5/20  ???  Y very little  H  18  5  85  G  Y  Y  Y  6.25  1/2  65-75  285/290  2085.0  G  some  check up, norhof offline for some reason, will defrost and restart
 2017/5/21 19:00  F  Y 50ml  H  18  5  85  G  Y  Y  Y  6.5  1/2  65-75  285-290  2109.0  G  some  filled norhof
 2017/5/22 20:12  30.1  Y50ML  H  18  5  85  G  Y  Y  Y  6.5  3/4  65-75  285-290  2134.1  G  some  checkup
 2017/5/22 07:45  29.4  Y 25 ml  H  N/A  N/A  85  G  Y  Y  Y  10.25  F  210  205  2143.8  G  some  shutdown
 2017/5/25 16:23  26.7  N  H  N/A  N/A  0  G  N  N  N  10.25  3/4  210  205  2143.8  G  some  changed oil in air compressor(6 MOS. interval)
 2017/6/8  3  Y  H      85  Y  Y  Y  Y  0  empty  210  205    G-Y  some  Startup by Greg (Orfeo assist). Dewar was empty so we purged for an hour then started He compressor. Startup went well. Some daily checks weren't recorded. 
 2017/6/10  17 ?                      1/2?            filled Norhof, unsure on LN2 levels
 2017/6/17  17?                    10.4 7/8             ln2 check, left running.
2017/9/13 - - - - - - - - - - - - - - - - -

While Greg did his daily checks he noticed the compressor was extremely hot and the piping leaving the intake filter was black from heat (blue quincy compressor head also discoloured) . When our operator turned the compressor off he felt very warm air leaving the inlet valve indicating it wasn't closing correctly. From all of this we suspect carbon has built up on the valve, it hasn't been able to close and the pump had to work hard to reach the shutoff at 110 PSI. 

Orfeo is post dating this on 18/6/2018 as i look for failure history on the QR-25. After this overheating failure i'm quite sure we ordered some QR-25 replacement parts and Greg rebuilt the compressor. Looking through emails we ordered MT-2074 UTO.pdf from Max Touckly <mtouckly@ncaltd.ca>

 

2017/10/12                                   Cryomech over voltage protection cutting power to He compressor. Voltage is 243 VAC.
 2017/10/13  0  y    18    85  y  y  y  y  4 psi  0  55-58  270-275        Cryomech over voltage protection cut power in the morning but then allowed power in the afternoon. Voltage 245 VAC at plug and electrical protection device. Helium pressure (off) 210,205. Voltage of legs of 240AC are 122,123. 245 VAC between A1/A2 of ABB voltage regulator. Filter bank filters replaced. Carbon filter was loose inside housing.
 2018/2/6                                    Greg has mentioned that the Quincy QR-25 air compressor won't run. He mentioned to Debra that the LN2 compressor will not stay on. It will blow the breaker and need to be rotated a bit by hand before it will restart. The last time he tried to fill the LN2 dewar, it seems to have created a vacuum, presumably since the generator was on, but the compressor was not. 

Greg thinks that this began after the lab temperatures dropped to ~-15C. 

 2018/03/05            87  G  Y  Y  Y  7-8  0 40  250  3518.5      Kris Kunz has started to take over operations of Eagle Bay Resort and the lab. I've asked him to run the LN2 system with Greg's old compressor. With He compressor off Low = 180 PSI, High=175. Kris purged dewar for an hour but pressure in dewar dropped to 0. Purged for 1 more hour and start up He compressor. 
 2018/03/06                      5              Dewar pressure trimmed up to 7 psi.
 2018/03/06                        1/4            checked in evening

 

Date & Time Norhof LN2 Level (cm)

Drain 

Air 
Comp.    

Air 

Comp.

Oil 

Lvl

Air

Comp.

Oil

Pres.

PSI

Air

Comp. 

Off

Period

Filter

Bank

Pres.

PSI

Water

Filter

 

ADV

Check

Sight

Carbon &

water

Filter

sight glass

Drain

carbon

filter

Dewar

Pres.

PSI

Dewar

LN2

Level

He

Low

Pres.

PSI

He

High

Pres.

PSI

He 

Comp.

Hours.

 

Air

Dryer

Dew
Point

Air

Dryer

water

Level

Notes:______
 YYYY/MM/DD  XX.X   Y/N  H/L  XX  X  XX  G/Y/R  Y/N  Y/N  Y/N  XX.X  X/4-8 XXX  XXX  XXXX.X  G/Y  X  

 

 The description of what's requested in the table is as follows:

 

Checklist: 

  1. Date & Time 
    1. Enter Date (YYYY/MM/DD) and Time (HH:MM in 24hr format)
  2. Drain Air Comp.  

    1. Drain Air Compressor manual drain valve on bottom of tank (Y=yes, N= No; volume)

  3. Air Comp. Oil Lvl
    1. Record Air Compressor Oil level several minutes after compressor has run (H = High, M= Medium, L= Low)
  4. Air Comp. Oil. Pres.
    1. Record Air Compressor Oil pressure (hot) while running, maintain 18-20 PSI. Adjust if necessary. 
  5. Air Comp. Off Period
    1. Record the period (minutes) the Air Compressor remains "off", time between "off" and next "on" cycle. Excessively long "off" periods can indicate cold head of Helium Compressor needs a service. At installation the "off" period was 6-7 min.
  6. Filter Bank Pres. 
    1.  Record Filter Bank regulator pressure (PSI)
  7. Water Filter.
    1. Record Water Filter Visual indicator, (G = green clean, Y=yellow , R= red change)
  8. ADV check
    1. Activate Automatic Drain Valve (ADV) 2 times and check that no water remains in the drain line after the second activation
  9. Sight Carbon and Water Filter sight glass
    1. Inspect sight glass / viewing windows of Carbon Filter and Water filter to make sure no water is present ( Y = yes done, N = No not done)
  10. Drain Carbon Filter  
    1. Y = yes done, N = No not done
  11. Dewar Pres.
    1. Record Dewar Pressure (PSI). Pressure should be 3-7 PSI (operating) and 10-12 PSI (compressors off)
    2.  If pressures greater than 15 PSI disconnect nitrogen flex line and open dewar bleed valve 
  12. Dewar LN2 Level
    1. Record the Liquid Nitrogen Level in the Dewar, fractional estimate is fine. 
  13. He Low Pres.
    1. Record Helium Compressor Low pressure gauge (PSI)
    2. Low gauge should read 210 PSI (off) or 70-80 PSI (operating)
  14. He High Pres.
    1. Record Helium Compressor High pressure gauge (PSI)
    2. High gauge should read 205 PSI (off) or 280-285 PSI (operating)
  15. He Comp. Hours
    1. Record Helium Compressor Hour meter value
  16. Air Dryer Dew Point
    1. Record Air Dryer Dew point indicator (G = Green, B = Blue, R= Red)
  17. Air Dryer water level
    1. Check water level in air dryer drain bottle. (Y= yes done, N= No not done)

 

 

Tag page